# On the conjugates of Christoffel words

### v3

@@ 1. Introduction
3. They constitute the finitary version of the Sturmian (infinite) words, which are obtained by discretizing straight lines in the plane, and which are characterized by the property that for each $n$, they have exactly $n+1$ factors of length $n$. For example, a word is a conjugate of a Christoffel word if and only if all its conjugates are factors of a Sturmian word; equivalently, this word (of length $n$ say) is primitive and has exactly $n-1$ circular factors of length $n-2$; see [L], [Re, Theorem 15.3.1].
@@ 2. Continuant polynomials
$[n_{1},\ldots,n_{k}]=n_{1}+\frac{1}{n_{2}+\frac{1}{n_{3}+\cdots+\frac{1}{n_{k}}}}$
is equal to the reduced fraction $K(n_{1},\ldots,n_{k})/K(n_{2},\ldots,n_{k})$. Equivalently, the continued fraction $[0,n_{1},\ldots,n_{k}]$ is equal to $K(n_{2},\ldots,n_{k})/K(n_{1},\ldots,n_{k})$.
@@ 3. Ostrowski numeration
For later use, we state the following result. We say that a sequence $d_{1},\ldots,d_{k}$ is alternating if its values are alternatively 00 and $b_{i}$; there are therefore two alternating sequences of length $k$.
@@ 6. Christoffel words
Equivalently, $S^{-1}=1+s^{-1}$. We have $S=0$ if and only if $s=0$, and $S=1$ if and only if $s=\infty$. Otherwise, $10<S<$ 1, and the continued fraction of $S$ is of the form $[0,a_{1},\ldots,a_{m}]$, where the $a_{i}$ are positive integers. Then $s^{-1}=S^{-1}-1=[a_{1},\ldots,a_{m}]-1=[a_{1}-1,a_{2},\ldots,a_{m}]$ if $a_{1}\geq 2$, and therefore $s=[0,a_{1}-1,a_{2},\ldots,a_{m}]$; and if $a_{1}=1$, we have $s^{-1}=[0,a_{2},\ldots,a_{m}]$ hence $s=[a_{2},\ldots,a_{m}]$.
@@ 7. Construction of the conjugates of a Christoffel word
Consider the alternating sequence beginning by 00, namely: $d_{1}=0,d_{2}=b_{2},\ldots$. Then $V_{1}=a^{b_{1}}b$ begins by $a$ if $a_{1}\geq 2$, and is equal to $b$ if $a_{1}=1$. Thus, if $a_{1}\geq 2$, then $V_{0},V_{1}$ begin by $a$, hence also do all $V_{i}$, $i\geq 0$. If $a_{1}=1$, then $V_{2}=V_{0}V_{1}^{b_{2}}=ab^{b_{2}}$, and $V_{3}=V_{2}^{b_{3}}V_{1}$ both begin by $a$; hence $V_{i}$ begins by $a$ for $i\geq 2$.
@@
Throughout the paper ${\widetilde{x}}$ denotes the reversal (mirror image) of the word $x$. A palindrome is a word equal to its reversal. The empty word is a palindrome. Recall that each proper lower Christoffel word $w$ has the factorization $w=apb$, where $p$ is a palindrome (called a central word), and that then the corresponding upper Christoffel word is $\widetilde{w}=bpa$, which is a conjugate of $w$. A standard word is $a$ or $b$, or a word of the from $pab$ or $pba$ for some central word $p$; it is known that standard words are obtained from $a,b$ by applying the endomorphisms in the submonoid generated by $G$ and $D$; moreover, each Christoffel class contains exactly two standard words $pab$ and $pba$, where $p$ is the corresponding central palindrome. See [L, Subsection 2.2.1], [Re].
@@
For $m\geq 1$, the word $M_{m}$ is a standard word, equal to $pab$ if $m$ is odd and to $pba$ if $m$ is even.
@@ 8.1. Borders
Suppose that $m\geq 3$ or $m=2$ and $b_{1}\geq 1$. Let $N$ be an integer with $10\leq N\leq q_{m}-$ 1 and $N=\sum_{1\leq i\leq m}d_{i}q_{i-1}$ be its greedy representation. Put $V_{m}=V_{m}(d_{1},\ldots,d_{m})$, $V_{m-1}=V_{m-1}(d_{1},\ldots,d_{m-1})$, and $V_{m-2}=V_{m-2}(d_{1},\ldots,d_{m-2})$. Assume that $V_{m}$ is not a Christoffel word. Let
@@
    If $11\leq d_{m}\leq b_{m}-$ 1 and $d_{m-1}=0$, then $B=V_{m-1}^{\ell}$, except if $b_{m}-d_{m}<d_{m}$ and the sequence $d_{1},\ldots,d_{m-1}$ is not alternating, in which case $B=V_{m-1}^{\ell+1}$.
@@
The key point for the proof of Theorem 8.1 is the determination of all the occurrences of $V_{m-1}$ in $V_{m}$. Exactly $b_{m}$ of them can be read on the factorisation (8.1), but there may be additional ones. By primitivity of $V_{m-1}$, the word $V_{m-1}V_{m-1}$ contains exactly two occurrences of $V_{m-1}$. Consequently, if an additional occurrence of $V_{m-1}$ appears, then it must be a factor of $V_{m-1}V_{m-2}V_{m-1}$. A more precise statement is given in Lemma 8.14.
@@ 8.2. Consequences
If $10\leq N<q_{m-1}-$ 1, then $B_{N}=H_{N}(a_{1},\ldots,a_{m-2})^{t}$, where $t=1+\min\{\lfloor\frac{N}{q_{m-2}}\rfloor,\lfloor\frac{q_{m-1}-2-N}{q_{m-2}}\rfloor\}$.
@@
We show that the numerator in the latter fraction is always in the interval $[-q_{m-1},q_{m-1})$, so that the integer part of this fraction is either --1 or 0, and we give the condition when it is 00. We have indeed $-q_{m-1}\leq q_{m-2}-2-j$, since $j\leq q_{m-1}-1$ by Lemma 10.2 (i), so that $j+2\leq q_{m-1}+1\leq q_{m-1}+q_{m-2}$. Moreover, $q_{m-2}-2-j<q_{m-1}$, since the sequence $q_{i}$ is increasing. Also, if $d_{m-1}>0$, then $j\geq q_{m-2}$, hence $q_{m-2}-2-j<0$; and if $d_{m-1}=0$, then $j=\sum_{1\leq i\leq m-2}d_{i}q_{i-1}$, and $q_{m-2}-2-j<0$ if and only if $j\geq q_{m-2}-1$, which, by Lemmas 10.1 (i) and 10.2 (i), is equivalent to the fact that the sequence $d_{1},\ldots,d_{m-2},0$ is alternating.
@@
4. We assume that $a_{m}\geq 2$. Suppose that $10\leq N<q_{m-1}-$ 1. Then $d_{m}=0$ by Lemma 10.2 (i). By Theorem 8.1 (v), we have $B_{N}=V_{m-1}^{b_{m}-1}V_{m-2}=H_{N}(a_{1},\ldots,a_{m-1},a_{m}-1)$.
@@
We have $H_{N}(a_{1},\ldots,a_{m})=C^{N}(M_{m})=C^{N}(V_{m}(0,\ldots,0))=C^{N}(V^{\prime}_{m-1}(0,\ldots,0,1))$ (by Lemma 8.5) $=C^{N}(C^{q_{m-2}}(M^{\prime}_{m-1}))$ (by Theorem 7.3 and because $q^{\prime}_{m-2}=q_{m-2}$) $=C^{N+q_{m-2}}(M^{\prime}_{m-1})=H_{N+q_{m-2}}(a^{\prime}_{1},\ldots,a^{\prime}_{m-1})$. Thus $H_{N}(a_{1},\ldots,a_{m})=H_{N^{\prime}}(a^{\prime}_{1},\ldots,a^{\prime}_{m-1})$ where $N^{\prime}=N+q_{m-2}$.
Suppose that $10\leq N<q_{m-1}-$ 1. Then $q^{\prime}_{m-2}=q_{m-2}\leq N^{\prime}<q_{m-1}+q_{m-2}-1=q_{m}-1=q^{\prime}_{m-1}-1$. It follows from Case (a) that $B_{N}=H_{N^{\prime}}(a^{\prime}_{1},\ldots,a^{\prime}_{m-2})^{t}$, with $t=\min\{\lfloor\frac{N^{\prime}}{q^{\prime}_{m-2}}\rfloor,1+\lfloor\frac{q^{\prime}_{m-1}-2-N^{\prime}}{q^{\prime}_{m-2}}\rfloor\}$. Note that $H_{N^{\prime}}(a^{\prime}_{1},\ldots,a^{\prime}_{m-2})=C^{N^{\prime}}(H_{0}(a^{\prime}_{1},\ldots,a^{\prime}_{m-2}))=C^{N+q_{m-2}}(H_{0}(a_{1},\ldots,a_{m-2}))=C^{N}(H_{0}(a_{1},\ldots,a_{m-2}))$ (since the word is of length $q_{m-2}$) $=H_{N}(a_{1},\ldots,a_{m-2})$; moreover, $t=\min\{\lfloor\frac{N+q_{m-2}}{q_{m-2}}\rfloor,1+\lfloor\frac{q_{m-1}+q_{m-2}-2-N-q_{m-2}}{q_{m-2}}\rfloor\}$, which settles this case.
Suppose now that $q_{m-1}\leq N<q_{m}-1$. Then $H_{N}(a_{1},\ldots,a_{m})=H_{N^{\prime}}(a^{\prime}_{1},\ldots,a^{\prime}_{m-1})=H_{N^{\prime\prime}}(a^{\prime}_{1},\ldots,a^{\prime}_{m-1})$, where $N^{\prime\prime}=N^{\prime}-q_{m}=N+q_{m-2}-q_{m}=N-q_{m-1}$, since the words have length $q_{m}$. Now $10\leq N^{\prime\prime}<q_{m-2}-$ 1. Hence by the first part, $B_{N}=H_{N^{\prime\prime}}(a^{\prime}_{1},\ldots,a^{\prime}_{m-1}-1)=C^{N^{\prime\prime}}(H_{0}(a_{1},\ldots,a_{m-1}))=C^{N}(H_{0}(a_{1},\ldots,a_{m-1}))$ (since the word has length $q_{m-1}$) $=H_{N}(a_{0},\ldots,a_{m-1})$. ∎
@@
This shows that the lemma holds for $k=0$. Now let $k\geq 0$ be an integer with $k<m-1$. Assume that $V_{k+1}V_{k}=W_{k}Z_{(-1)^{k+1}}X_{k}$ and $V_{k}V_{k+1}=W_{k}Z_{(-1)^{k}}X_{k}$.
@@
Remark. Observe that if $UU$ is a border of $W$ and $V$ a border of $U$, then $UV$ is not necessarily a border of $W$. A counterexample is given by $abaababbabaaba$, with $U=aba$ and $V=a$.
@@
1. Assume first that $11\leq d_{m}\leq b_{m}-$ 1. Thus by (8.1), the word $V_{m-1}$ is a prefix and a suffix of $V_{m}$. We show that either $X$ is an integer power of $V_{m-1}$, or $Y$ is an integer power of $V_{m-1}$. Indeed, if $|X|<|V_{m-1}|$, then $X$ is a nontrivial proper suffix of $V_{m-1}$, $V_{m-1}=UX$, where $U$ is nonempty, and $V_{m}$ begins with $XV_{m-1}$; but $V_{m}$ also begins with $V_{m-1}$, $V_{m}=V_{m-1}W$, hence $UV_{m-1}W=UV_{m}=UXV_{m-1}Y=V_{m-1}V_{m-1}Y$; since $U$ is nonempty and shorter than $V_{m-1}$, we see that $V_{m-1}$ is a proper factor of $V_{m-1}^{2}$, and we have a contradiction with the primitivity of $V_{m-1}$.
@@
1. Suppose by contradiction that $B=Y^{i}XY^{r}$, with $0\leq r\leq j$. Then $r<j$ since $B\neq W$. Moreover $Y^{i}XY^{j-r}Y^{r}=W=UB=UY^{i-1}YXY^{r}$, hence $Y^{i}XY^{j-r}=UY^{i-1}YX$; if $j-r=1$, then $XY=YX$, a contradiction; thus $j-r\geq 2$ and, since $X$ is shorter than $Y$, we see that $Y$ is an internal factor of $YY$, which contradicts the primitivity of $Y$.
@@
We have $d_{m-1}=0$. Suppose that the lemma is false. We show first that $V_{m}$ is a Christoffel word. It is enough to prove that the $d_{i}$ are alternatively $b_{i}$ and 0 (Corollary 7.4). Since $V_{m}=V_{m-2}V_{m-1}^{b_{m}}$ is a suffix of $V_{m}V_{m-1}=V_{m-2}V_{m-1}^{b_{m}+1}$, the word $V_{m-2}$ is a suffix of $V_{m-1}$. Note that for $m=1$, this cannot be true, and neither for $m=2$, since $V_{1}=a^{b_{1}}b$, $V_{0}=a$; hence we must have $m\geq 3$. Since $d_{m-1}=0$, we deduce from Lemma 8.17, applied to $m-1$, that $d_{m-2}=b_{m-2}$.
@@
Observe that $V_{m-1}=V_{m-2}^{b_{m-1}}V_{m-3}$ is a prefix of $V_{m}=V_{m-2}V_{m-2}^{b_{m-1}}V_{m-3}\cdots$ if and only if $V_{m-3}$ is a prefix of $V_{m-2}$, thus, by Corollary 8.9, if and only if $V_{m-2}$ is not the Christoffel word $V_{m-2}(\ldots,0,b_{m-2})$. But $V_{m-2}$ cannot be equal to the latter word, since $V_{m}$ is by assumption not a Christoffel word, and $d_{m}=b_{m},d_{m-1}=0$. Thus $V_{m-1}$ is a prefix of $V_{m}$, and $B=V_{m-1}$ by Lemmas 8.13 (applied to $m-1$) and 8.16.
$(ii)$ Suppose that $11\leq d_{m}\leq b_{m}-$ 1 and $11\leq d_{m-1}\leq b_{m-1}-$ 1. There are no other occurrences of $V_{m-1}$ in $V_{m}$ than those seen in the factorization $V_{m}=V_{m-1}^{b_{m}-d_{m}}V_{m-2}V_{m-1}^{d_{m}}$; indeed, this follows from Lemma 8.14 (applied to $m-1$), and our assumption on $d_{m-1}$. Consequently, by Lemma 8.15, $B=V_{m-1}^{\ell}$.
$(iii)$ If $11\leq d_{m}\leq b_{m}-$ 1 and $d_{m-1}=0$, then $V_{m-1}=V_{m-2}^{b_{m-1}}V_{m-3}$. By Lemma 8.14 (applied to $m-1$), and the hypothesis $d_{m-1}=0$, any occurrence of $V_{m-1}$ in $V_{m}$ can be read on the factorisation $V_{m}=V_{m-1}^{b_{m}-d_{m}}V_{m-2}V_{m-1}^{d_{m}}$, or it begins by $V_{m-2}$. It follows from Lemma 8.15 that $B=V_{m-1}^{\ell+1}$ if $V_{m-1}$ is a prefix of $V_{m-2}V_{m-1}$ and $b_{m}-d_{m}<d_{m}$, and otherwise $B=V_{m-1}^{\ell}$.
@@
$(iv)$ Suppose that $11\leq d_{m}\leq b_{m}-$ 1 and $d_{m-1}=b_{m-1}$. Then $V_{m-1}=V_{m-3}V_{m-2}^{b_{m-1}}$.
@@
Supppose now that $V_{m-1}$ is a suffix of $V_{m-1}V_{m-2}$. This contradicts Lemma 8.18, applied to $m-1$, since $d_{m-2}=0$ by greedyness.
@@
If $b_{m-1}-d_{m-1}\geq 1$ and $11\leq d_{m-2}\leq b_{m-2}-$ 1, then it follows from Lemma 8.14 (applied to $m-2$) and the hypothesis on $d_{m-2}$, that there are no further occurrences of $V_{m-2}$ in $V_{m}$. Thus $B=V_{m-2}^{h}$ by Lemma 8.15.
If $b_{m-1}-d_{m-1}\geq 1$ and $d_{m-2}=0$, then by Lemmas 8.14 and 8.15, $B=V_{m-2}^{h+1}$ if $b_{m-1}-d_{m-1}<d_{m-1}+1$ and $V_{m-2}$ is a prefix of $V_{m-3}V_{m-2}$, and $B=V_{m-2}^{h}$ otherwise. But, by Corollary 8.10 with $m$ replaced by $m-1$, $V_{m-2}$ is a prefix of $V_{m-3}V_{m-2}$ if and only if the sequence $d_{1},\ldots,d_{m-2}$ is not alternating.
If $b_{m-1}-d_{m-1}\geq 1$ and $d_{m-2}=b_{m-2}$, then by Lemma 8.18 with $m$ replaced by $m-2$, $V_{m-2}$ is not a suffix of $V_{m-2}V_{m-3}$. Thus by Lemmas 8.14 and 8.15, $B=V_{m-2}^{h}$.

### v4

@@ 1. Introduction
3. They constitute the finitary version of the Sturmian (infinite) words, which are obtained by discretizing straight lines in the plane, and which are characterized by the property that for each $n$, they have exactly $n+1$ factors of length $n$. For example, a word is a conjugate of a Christoffel word if and only if all its conjugates are factors of a Sturmian word; equivalently, this word (of length $n$ say) is primitive and has exactly n1n-1 circular factors of length n2n-2; see [L], [Re, Theorem 15.3.1].
@@ 2. Continuant polynomials
$n_{1},\ldots,n_{k}]=n_{1}+\frac{1}{n_{2}+\frac{1}{n_{3}+\cdots+\frac{1}{n_{k}}}}$
is equal to the reduced fraction $K(n_{1},\ldots,n_{k})/K(n_{2},\ldots,n_{k})$. Equivalently, the continued fraction $0,n_{1},\ldots,n_{k}]$ is equal to $K(n_{2},\ldots,n_{k})/K(n_{1},\ldots,n_{k})$.
@@ 3. Ostrowski numeration
For later use, we state the following result. We say that a sequence $d_{1},\ldots,d_{k}$ is alternating if its values are alternatively 0 and $b_{i}$; there are therefore two alternating sequences of length $k$.
@@ 6. Christoffel words
Equivalently, $S^{-1}=1+s^{-1}$. We have $S=0$ if and only if $s=0$, and $S=1$ if and only if $s=\infty$. Otherwise, $0<S<1$, and the continued fraction of $S$ is of the form $0,a_{1},\ldots,a_{m}]$, where the $a_{i}$ are positive integers. Then $s^{-1}=S^{-1}-1=[a_{1},\ldots,a_{m}]-1=[a_{1}-1,a_{2},\ldots,a_{m}]$ if $a_{1}\geq 2$, and therefore $s=[0,a_{1}-1,a_{2},\ldots,a_{m}]$; and if $a_{1}=1$, we have $s^{-1}=[0,a_{2},\ldots,a_{m}]$ hence $s=[a_{2},\ldots,a_{m}]$.
@@ 7. Construction of the conjugates of a Christoffel word
Consider the alternating sequence beginning by 0, namely: $d_{1}=0,d_{2}=b_{2},\ldots$. Then $V_{1}=a^{b_{1}}b$ begins by $a$ if $a_{1}\geq 2$, and is equal to $b$ if $a_{1}=1$. Thus, if $a_{1}\geq 2$, then $V_{0},V_{1}$ begin by $a$, hence also do all $V_{i}$, $i\geq 0$. If $a_{1}=1$, then $V_{2}=V_{0}V_{1}^{b_{2}}=ab^{b_{2}}$, and $V_{3}=V_{2}^{b_{3}}V_{1}$ both begin by $a$; hence $V_{i}$ begins by $a$ for $i\geq 2$.
@@
Throughout the paper $\widetilde{x}$} denotes the reversal (mirror image) of the word $x$. A palindrome is a word equal to its reversal. The empty word is a palindrome. Recall that each proper lower Christoffel word $w$ has the factorization $w=apb$, where $p$ is a palindrome (called a central word), and that then the corresponding upper Christoffel word is $\widetilde{w}=bpa$, which is a conjugate of $w$. A standard word is $a$ or $b$, or a word of the from pabpab or pbapba for some central word $p$; it is known that standard words are obtained from $a,b$ by applying the endomorphisms in the submonoid generated by $G$ and $D$; moreover, each Christoffel class contains exactly two standard words pabpab and pbapba, where $p$ is the corresponding central palindrome. See [L, Subsection 2.2.1], [Re].
@@
For $m\geq 1$, the word $M_{m}$ is a standard word, equal to pabpab if $m$ is odd and to pbapba if $m$ is even.
@@ 8.1. Borders
Suppose that $m\geq 3$ or $m=2$ and $b_{1}\geq 1$. Let $N$ be an integer with $0\leq N\leq q_{m}-1$ and $N=\sum_{1\leq i\leq m}d_{i}q_{i-1}$ be its greedy representation. Put $V_{m}=V_{m}(d_{1},\ldots,d_{m})$, $V_{m-1}=V_{m-1}(d_{1},\ldots,d_{m-1})$, and $V_{m-2}=V_{m-2}(d_{1},\ldots,d_{m-2})$. Assume that $V_{m}$ is not a Christoffel word. Let
@@
    If $1\leq d_{m}\leq b_{m}-1$ and $d_{m-1}=0$, then $B=V_{m-1}^{\ell}$, except if $b_{m}-d_{m}<d_{m}$ and the sequence $d_{1},\ldots,d_{m-1}$ is not alternating, in which case $B=V_{m-1}^{\ell+1}$.
@@
The key point for the proof of Theorem 8.1 is the determination of all the occurrences of $V_{m-1}$ in $V_{m}$. Exactly $b_{m}$ of them can be read on the factorization (8.1), but there may be additional ones. By primitivity of $V_{m-1}$, the word $V_{m-1}V_{m-1}$ contains exactly two occurrences of $V_{m-1}$. Consequently, if an additional occurrence of $V_{m-1}$ appears, then it must be a factor of $V_{m-1}V_{m-2}V_{m-1}$. A more precise statement is given in Lemma 8.14.
@@ 8.2. Consequences
As noted by one of the referees, this result may also be obtained as follows: if $u$ is a border of the Christoffel word $w$, then $uu$ is a factor of $ww$; since $ww$ is a Sturmian word, it is Sturmian, and thus $uu$ too. Hence all conjugates of $u$ are Sturmian. Among them is the power $\ell^{k}$ of some Lyndon word $\ell$, which is therefore Sturmian; hence $\ell$ is a Christoffel word by a theorem of Berstel and de Luca (see [Re] Corollary 13.4.3).
@@
If $0\leq N<q_{m-1}-1$, then $B_{N}=H_{N}(a_{1},\ldots,a_{m-2})^{t}$, where $t=1+\min\{\lfloor\frac{N}{q_{m-2}}\rfloor,\lfloor\frac{q_{m-1}-2-N}{q_{m-2}}\rfloor\}$.
@@
We show that the numerator in the latter fraction is always in the interval $q_{m-1},q_{m-1})$, so that the integer part of this fraction is either -1 or 0, and we give the condition when it is 0. We have indeed $q_{m-1}\leq q_{m-2}-2-j$, since $j\leq q_{m-1}-1$ by Lemma 10.2 (i), so that $j+2\leq q_{m-1}+1\leq q_{m-1}+q_{m-2}$. Moreover, $q_{m-2}-2-j<q_{m-1}$, since the sequence $q_{i}$ is increasing. Also, if $d_{m-1}>0$, then $j\geq q_{m-2}$, hence $q_{m-2}-2-j<0$; and if $d_{m-1}=0$, then $j=\sum_{1\leq i\leq m-2}d_{i}q_{i-1}$, and $q_{m-2}-2-j<0$ if and only if $j\geq q_{m-2}-1$, which, by Lemmas 10.1 (i) and 10.2 (i), is equivalent to the fact that the sequence $d_{1},\ldots,d_{m-2},0$ is alternating.
@@
4. We assume that $a_{m}\geq 2$. Suppose that $0\leq N<q_{m-1}-1$. Then $d_{m}=0$ by Lemma 10.2 (i). By Theorem 8.1 (v), we have $B_{N}=V_{m-1}^{b_{m}-1}V_{m-2}=H_{N}(a_{1},\ldots,a_{m-1},a_{m}-1)$.
@@
We have $H_{N}(a_{1},\ldots,a_{m})=C^{N}(M_{m})=C^{N}(V_{m}(0,\ldots,0))=C^{N}(V^{\prime}_{m-1}(0,\ldots,0,1))$ (by Lemma 8.5) $C^{N}(C^{q_{m-2}}(M^{\prime}_{m-1}))$ (by Theorem 7.3 and because $q^{\prime}_{m-2}=q_{m-2}$) $C^{N+q_{m-2}}(M^{\prime}_{m-1})=H_{N+q_{m-2}}(a^{\prime}_{1},\ldots,a^{\prime}_{m-1})$. Thus $H_{N}(a_{1},\ldots,a_{m})=H_{N^{\prime}}(a^{\prime}_{1},\ldots,a^{\prime}_{m-1})$ where $N^{\prime}=N+q_{m-2}$.
Suppose that $0\leq N<q_{m-1}-1$. Then $q^{\prime}_{m-2}=q_{m-2}\leq N^{\prime}<q_{m-1}+q_{m-2}-1=q_{m}-1=q^{\prime}_{m-1}-1$. It follows from Case (a) that $B_{N}=H_{N^{\prime}}(a^{\prime}_{1},\ldots,a^{\prime}_{m-2})^{t}$, with $t=\min\{\lfloor\frac{N^{\prime}}{q^{\prime}_{m-2}}\rfloor,1+\lfloor\frac{q^{\prime}_{m-1}-2-N^{\prime}}{q^{\prime}_{m-2}}\rfloor\}$. Note that $H_{N^{\prime}}(a^{\prime}_{1},\ldots,a^{\prime}_{m-2})=C^{N^{\prime}}(H_{0}(a^{\prime}_{1},\ldots,a^{\prime}_{m-2}))=C^{N+q_{m-2}}(H_{0}(a_{1},\ldots,a_{m-2}))=C^{N}(H_{0}(a_{1},\ldots,a_{m-2}))$ (since the word is of length $q_{m-2}$) $H_{N}(a_{1},\ldots,a_{m-2})$; moreover, $t=\min\{\lfloor\frac{N+q_{m-2}}{q_{m-2}}\rfloor,1+\lfloor\frac{q_{m-1}+q_{m-2}-2-N-q_{m-2}}{q_{m-2}}\rfloor\}$, which settles this case.
Suppose now that $q_{m-1}\leq N<q_{m}-1$. Then $H_{N}(a_{1},\ldots,a_{m})=H_{N^{\prime}}(a^{\prime}_{1},\ldots,a^{\prime}_{m-1})=H_{N^{\prime\prime}}(a^{\prime}_{1},\ldots,a^{\prime}_{m-1})$, where $N^{\prime\prime}=N^{\prime}-q_{m}=N+q_{m-2}-q_{m}=N-q_{m-1}$, since the words have length $q_{m}$. Now $0\leq N^{\prime\prime}<q_{m-2}-1$. Hence by the first part, $B_{N}=H_{N^{\prime\prime}}(a^{\prime}_{1},\ldots,a^{\prime}_{m-1}-1)=C^{N^{\prime\prime}}(H_{0}(a_{1},\ldots,a_{m-1}))=C^{N}(H_{0}(a_{1},\ldots,a_{m-1}))$ (since the word has length $q_{m-1}$) $H_{N}(a_{0},\ldots,a_{m-1})$. ∎
@@
This shows that the lemma holds for $k=0$. Now let $k\geq 0$ be an integer with k<m1k<m-1. Assume that $V_{k+1}V_{k}=W_{k}Z_{(-1)^{k+1}}X_{k}$ and $V_{k}V_{k+1}=W_{k}Z_{(-1)^{k}}X_{k}$.
@@
Remark. Observe that if $UU$ is a border of $W$ and $V$ a border of $U$, then $UV$ is not necessarily a border of $W$. A counterexample is given by abaababbabaabaabaababbabaaba, with $U=aba$ and $V=a$.
@@
1. Assume first that $1\leq d_{m}\leq b_{m}-1$. Thus by (8.1), the word $V_{m-1}$ is a prefix and a suffix of $V_{m}$. We show that either $X$ is an integer power of $V_{m-1}$, or $Y$ is an integer power of $V_{m-1}$. Indeed, if $|X|<|V_{m-1}|$, then $X$ is a nontrivial proper suffix of $V_{m-1}$, $V_{m-1}=UX$, where $U$ is nonempty, and $V_{m}$ begins with $XV_{m-1}$; but $V_{m}$ also begins with $V_{m-1}$, $V_{m}=V_{m-1}W$, hence $UV_{m-1}W=UV_{m}=UXV_{m-1}Y=V_{m-1}V_{m-1}Y$; since $U$ is nonempty and shorter than $V_{m-1}$, we see that $V_{m-1}$ is a proper factor of $V_{m-1}^{2}$, and we have a contradiction with the primitivity of $V_{m-1}$.
@@
1. Suppose by contradiction that $B=Y^{i}XY^{r}$, with $0\leq r\leq j$. Then $r<j$ since $B\neq W$. Moreover $Y^{i}XY^{j-r}Y^{r}=W=UB=UY^{i-1}YXY^{r}$, hence $Y^{i}XY^{j-r}=UY^{i-1}YX$; if jr=1j-r=1, then $XY=YX$, a contradiction; thus $j-r\geq 2$ and, since $X$ is shorter than $Y$, we see that $Y$ is an internal factor of $YY$, which contradicts the primitivity of $Y$.
@@
We have $d_{m-1}=0$. Suppose that the lemma is false. We show first that $V_{m}$ is a Christoffel word. It is enough to prove that the $d_{i}$ are alternatively $b_{i}$ and 0 (Corollary 7.4). Since $V_{m}=V_{m-2}V_{m-1}^{b_{m}}$ is a suffix of $V_{m}V_{m-1}=V_{m-2}V_{m-1}^{b_{m}+1}$, the word $V_{m-2}$ is a suffix of $V_{m-1}$. Note that for $m=1$, this cannot be true, and neither for $m=2$, since $V_{1}=a^{b_{1}}b$, $V_{0}=a$; hence we must have $m\geq 3$. Since $d_{m-1}=0$, we deduce from Lemma 8.17, applied to m1m-1, that $d_{m-2}=b_{m-2}$.
@@
Observe that $V_{m-1}=V_{m-2}^{b_{m-1}}V_{m-3}$ is a prefix of $V_{m}=V_{m-2}V_{m-2}^{b_{m-1}}V_{m-3}\cdots$ if and only if $V_{m-3}$ is a prefix of $V_{m-2}$, thus, by Corollary 8.9, if and only if $V_{m-2}$ is not the Christoffel word $V_{m-2}(\ldots,0,b_{m-2})$. But $V_{m-2}$ cannot be equal to the latter word, since $V_{m}$ is by assumption not a Christoffel word, and $d_{m}=b_{m},d_{m-1}=0$. Thus $V_{m-1}$ is a prefix of $V_{m}$, and $B=V_{m-1}$ by Lemmas 8.13 (applied to m1m-1) and 8.16.
$(ii)$ Suppose that $1\leq d_{m}\leq b_{m}-1$ and $1\leq d_{m-1}\leq b_{m-1}-1$. There are no other occurrences of $V_{m-1}$ in $V_{m}$ than those seen in the factorization $V_{m}=V_{m-1}^{b_{m}-d_{m}}V_{m-2}V_{m-1}^{d_{m}}$; indeed, this follows from Lemma 8.14 (applied to m1m-1), and our assumption on $d_{m-1}$. Consequently, by Lemma 8.15, $B=V_{m-1}^{\ell}$.
$(iii)$ If $1\leq d_{m}\leq b_{m}-1$ and $d_{m-1}=0$, then $V_{m-1}=V_{m-2}^{b_{m-1}}V_{m-3}$. By Lemma 8.14 (applied to m1m-1), and the hypothesis $d_{m-1}=0$, any occurrence of $V_{m-1}$ in $V_{m}$ can be read on the factorisation $V_{m}=V_{m-1}^{b_{m}-d_{m}}V_{m-2}V_{m-1}^{d_{m}}$, or it begins by $V_{m-2}$. It follows from Lemma 8.15 that $B=V_{m-1}^{\ell+1}$ if $V_{m-1}$ is a prefix of $V_{m-2}V_{m-1}$ and $b_{m}-d_{m}<d_{m}$, and otherwise $B=V_{m-1}^{\ell}$.
@@
$(iv)$ Suppose that $1\leq d_{m}\leq b_{m}-1$ and $d_{m-1}=b_{m-1}$. Then $V_{m-1}=V_{m-3}V_{m-2}^{b_{m-1}}$.
@@
Supppose now that $V_{m-1}$ is a suffix of $V_{m-1}V_{m-2}$. This contradicts Lemma 8.18, applied to m1m-1, since $d_{m-2}=0$ by greedyness.
@@
If $b_{m-1}-d_{m-1}\geq 1$ and $1\leq d_{m-2}\leq b_{m-2}-1$, then it follows from Lemma 8.14 (applied to m2m-2) and the hypothesis on $d_{m-2}$, that there are no further occurrences of $V_{m-2}$ in $V_{m}$. Thus $B=V_{m-2}^{h}$ by Lemma 8.15.
If $b_{m-1}-d_{m-1}\geq 1$ and $d_{m-2}=0$, then by Lemmas 8.14 and 8.15, $B=V_{m-2}^{h+1}$ if $b_{m-1}-d_{m-1}<d_{m-1}+1$ and $V_{m-2}$ is a prefix of $V_{m-3}V_{m-2}$, and $B=V_{m-2}^{h}$ otherwise. But, by Corollary 8.10 with $m$ replaced by m1m-1, $V_{m-2}$ is a prefix of $V_{m-3}V_{m-2}$ if and only if the sequence $d_{1},\ldots,d_{m-2}$ is not alternating.
If $b_{m-1}-d_{m-1}\geq 1$ and $d_{m-2}=b_{m-2}$, then by Lemma 8.18 with $m$ replaced by m2m-2, $V_{m-2}$ is not a suffix of $V_{m-2}V_{m-3}$. Thus by Lemmas 8.14 and 8.15, $B=V_{m-2}^{h}$.